OpenStudy (anonymous):

WILL FAN AND MEDAL What is the solution to the system of equations? x=3y+5 x-4y=3 ( , )

OpenStudy (trojanpoem):

x= 4y +3 4y +3 = 3y +5 2y = 2 y = 1 x = 4 +3 = 7 (7,1)